A machine is available 10 hours a day. Each part takes 90 minutes to fabricate and 10 minutes to setup 10% o the parts made are defect e and must go through the machine again to be reworked, during which both setup and fabrication have to be performed again. The machine is idle for any unused time Or average, 5 good parts are made each day. What is the overall equipment effectiveness of the machine? 
  • 0.17
  • 0.75
  • 0.075
  • 0.83
Answer #1

Time available each day \(=10\) hour

\(=10 \times 60\)

\(=600\) minutes

Time take to fabricate 1 part \(=90\) minutes

Time setup 1 part \(=10\) minutes

Total parts made in one day \(=5\)

Parts defective \(=10 \%\)

Total time to produce 1 part = time to setup + time to fabricate

\(=10+90\)

\(=100\) minutes

Thus, time taken to produce 5 parts \(=5 \times 100\)

\(=500\) minutes

Number of parts defective \(=10 \%\) of total parts made in one day

\(=0.10 \times 5\)

\(=0.5\) parts

Time taken to replace defective part \(=0.5 \times 100\)

\(=50\) minutes

Value added time = Time taken to produce 5 parts - time taken to replace defective part

\(=500-50\)

\(=450\) mimutes

Calculate effectiveness using following formula:

\(\begin{aligned} \text { Effectiveness } &=\frac{\text { value added time }}{\text { total time }} \\ &=\frac{450}{600} \\ &=0.75 \end{aligned}\)

Hence, correct answer is \(0.75\)